A women is five times as old as her grandson. If she is sixty years older than the boy, what are their ages?
Answer by ikleyn(52803)

.
    5x - x = 60,


as the problem says.


In other words,


    4x = 60  and  x = 60/4 = 15.


ANSWER.  The woman is  5*15 = 75 years old.  The grandson is 15 years old.